TEMPERATURE DEPENDENCE OF SELF- and NITROGEN-GAS LINE SHAPE PERTURBATIONS IN THE ν_{1} + ν_{3} BAND OF ACETYLENE

MATTHEW J. CICH, GARY V. LOPEZ, \underline{TREVOR J. SEARS} ^{a}, Department of Chemistry, Stony Brook University, Stony Brook, New York 11794; C. P. MCRAVEN, Department of Chemistry, Brookhaven National Laboratory, Upton, New York 11973; A. W. MANTZ, Department of Physics, Astronomy, and Astrophysics, Connecticut College, New London, CT 06320; and DANIEL HURT-MANS, Service de Chimie Quantique et de Photophysique(Atoms, Molecules et Atmospheres), Universite Libre de Bruxelles, Bruxelles, Belgium B-10050.

The calibration of remote sensing measurements of the atmosphere of our own and other planetary bodies requires a knowledge of the temperature dependence of spectral line broadening caused by pressure dependent effects. We have extended the frequency comb-referenced measurements of line broadening in the ν_{1} + ν_{3} band of acetylene at 296 K, reported elsewhere at this symposium, to include low temperature pressure measurements at 175K, 200K and 240K. Additional data at still lower temperatures is currently being obtained. Line shape parameters for self- and nitrogenbroadening, narrowing and shifts have been determined at these temperatures on the assumption of the hard-collision model. This model was found to be the most appropriate in our earlier 296K data fitting. In addition to its intrinsic value, the temperature dependence of the derived parameters can give insight into the reliability of the physical assumptions of the model, and results of our analysis will be presented at the meeting. Acknowledgements: Acknowledgement is made to the Donors of the American Chemical Society Petroleum Research Fund for partial support of this research. CPM gratefully acknowledges support by DOE EPSCoR grant DOE-07ER46361 for work conducted while at the University of Oklahoma. The measurements and analyses were performed under grants NNX09AJ93G and NNX08AO78G from the NASA Planetary and Atmospheres program.

aalso: Department of Chemistry, Brookhaven National Laboratory, Upton, New York 11973